OpenStudy (anonymous):

In the below system, solve for y in the first equation. x – 3y = –6 2x – 7y = 10

OpenStudy (anonymous):

need help really confused

OpenStudy (anonymous):

i just had something similiar to this earlier today let me see if i can try to help

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

do you have to do anything with the 2nd equation

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

y=6−x 3 y=6 3 −x 3 y=2−1 3 x y=−1 3 x+2 So the answer is choice D

OpenStudy (anonymous):

that's the answer

OpenStudy (anonymous):

i meant in general.. eventually you will take the 2 and combine them i think

OpenStudy (anonymous):

yeah I got it thanks for the help anyway

OpenStudy (anonymous):

you want the web site i was given to kind of help me

OpenStudy (anonymous):

www.wolframalpha.com